INFRARED CAVITY RINGDOWN LASER ABSORPTION SPECTROSCOPY

\underline{R. A. PROVENCAL}, J. B. PAUL, K. ROTH, R. CASAEAS, A. PETERSSON, and R. J. SAYKALLY, Department of Chemistry, UC Berkeley, Berkeley, CA 94720.

We are vigorously pursuing extension of cavity ringdown laser absorption spectroscopy into the infrared region.^{a}^{,}^{b} With medium resolution (˜ 1000 MHz), we have measured spectra of H_{2}O, D_{2}O, CH_{3}OH, CH_{3}CH_{2}OH, and CH_{3}(CH_{2})_{3}OH clusters in the spectral region between 2.8 and 6 microns. By employing a laser vaporization/supersonic molecular beam source, we have also studied carbon clusters in this spectral region. High resolution (<100 MHz) cavity ringdown experiments are also being pursued using a novel Alexandrite based laser system.

aJ. B. Paul, and R. J. Saykally. Anal. Chem. 69, 287A (1997).

bJ. B. Paul, R. A. Provencal, and R. J. Saykally. J. Phys. Chem. in press.